OpenStudy (anonymous):

Find the domain of the function f (x)=6-3x

OpenStudy (aroub):

6-3x=0 -3x=-6 x=2

OpenStudy (sriram):

all real numbers

OpenStudy (sriram):

domain is the values x is allowed to attain

OpenStudy (anonymous):

Domain is R.